OpenStudy (amorfide):

The random variable W has the distribution B(10,1/3). Find (a) P(W≤2) (b) P(W=2) (ii) The random variable X has the distribution B(15, 0.22) (a) Find P(X=4) (b) Find E(X) and Var(X)

OpenStudy (anonymous):

this a binomial probability right?

OpenStudy (amorfide):

yes

OpenStudy (anonymous):

so if i am not mistake \[P(X=k)=\dbinom{n}{k}p^k(1-p)^{n-k}\] in your case you have \[P(X=2)=\dbinom{10}{2}(\frac{1}{3})^2(\frac{2}{3})^8\]

OpenStudy (anonymous):

need a calculator for this. but \[\dbinom{10}{2}=\frac{10\times 9}{2}=45\]

OpenStudy (anonymous):

\[P(X\leq 2)=P(X=0)+P(X=1)+P(X=2)\] \[=\left(\frac{2}{3}\right)^{10}+10\left(\frac{1}{3}\right)\left(\frac{2}{3}\right)^9+\dbinom{10}{2}\left(\frac{1}{3}\right)^2\left(\frac{2}{3}\right)^8\]

OpenStudy (amorfide):

thank you very much! can't wait to get stats over and done with

OpenStudy (anonymous):

i am sure. but these are not that bad, got a formula and a calculator and you should be good to go

OpenStudy (amorfide):

it was more understanding how to go about the question... stats is my worst subject im better at the more complicated stuff